OpenStudy (anonymous):

At his trucking company, Richard received an order to ship 157 cubic feet of oil in cylindrical steel barrels. If each barrel has a radius of 1.1 feet and is 3.7 feet tall, how many steel barrels will Richard need?

OpenStudy (anonymous):

pi r^2 h = volume of single barrel 157/ volume of a single barrel = amount of barrels

OpenStudy (anonymous):

r=1.1 h=3.7

OpenStudy (anonymous):

So it's 638.99?

OpenStudy (anonymous):

Im so confused. My answer choices are 10 11 12 and 13.

OpenStudy (anonymous):

1.1*1.1*3.14*3.7=2207 and take the cube root so that would be about 13 but a little over/

OpenStudy (anonymous):

Essentially the equation is this 157/(pi *1.1 ^(2) * 3.7)

OpenStudy (anonymous):

i don't have my calc on me but no way does basically 1 *9 = 2207